MOA-II Obsevation Log Monthly Summary 2007/11

Return to index

Previous month Next month

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB11383-114335125115060
LMC1897-21052091080314100
SMC457-478224312720
ALERT-000
GRB-000
TOTAL---2821374341880 (94h58m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
121520829713103572411
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
7641064532130


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
434443383550190186180575152
Field13Field14Field15Field16
555600


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
22210000000000


SPECIAL TARGET NAME

CategoryTargetNrun


CCD TEMPERATURE Previous month Next month




VACUUM Previous month Next month




TELESCOPE TEMPERATURE Previous month Next month




COMMENTS
2007/11/1: DARK D2902(300sec), D2903(60sec) are OK. Pumping off / night time = 27000(sec)

2007/11/2: DARK D2904(300sec), D2905(60sec) are OK. / night time = 26760(sec)

2007/11/3: DARK D2907(300sec) is OK. Pumping on going / night time = 26520(sec)

2007/11/4: No observation. We took FLAT images(F2028-F2063). DARK D2910(15sec) is OK. Pumping off. / night time = 26280(sec)

2007/11/5: I could't take good DARKs, because chip2 was noisy. (D2936(60sec), D2937(300sec) for the record) / night time = 26040(sec)

2007/11/6: DARK D2976(300sec), D2977(60sec) are OK. / night time = 25800(sec)

2007/11/7: DARK D2983(300sec), D2984(60sec) are OK. / night time = 25560(sec)

2007/11/8: DARK D2988(300sec), D2989(60sec) are OK. / night time = 25320(sec)

2007/11/9: DARK D3046(300sec), D3047(60sec) are OK. / night time = 25020(sec)

2007/11/10: DARK D3057(300sec) is OK. / night time = 24780(sec)

2007/11/11: DARK D3066(300sec) is OK. / night time = 24600(sec)

2007/11/12: DARK D3073(300sec) is OK. / night time = 24360(sec)

2007/11/13: No observation / night time = 24120(sec)

2007/11/14: DARK D3108(300sec) is OK. / night time = 23880(sec)

2007/11/15: DARK D3113(300sec) is OK. / night time = 23640(sec)

2007/11/16: DARK D3128(300sec) is OK. / night time = 23400(sec)

2007/11/17: DARK D3137(300sec) is OK. Pumping on going. / night time = 23160(sec)

2007/11/18: DARK D3140(300sec) is OK. Pumping off / night time = 22920(sec)

2007/11/19: Change exposure time 240. DARK D3148(240sec), D3150(300sec) are OK. But D3150 is slightly noisy. / night time = 22680(sec)

2007/11/20: Pumping before observation, DARK D3155(300sec) is OK (it have some noise because of high CCD temperature), pumping on going / night time = 22440(sec)

2007/11/21: DARK D3161(300sec) is OK. Pumping off / night time = 22260(sec)

2007/11/22: change exposure time 240s 180s 120s, DARK D3168(240sec), D3166(180sec), D3165(120sec) are OK / night time = 22020(sec)

2007/11/23: I tried to change vacuum sensor. But I could not change. Pumping before observation, take FLAT(F2067-F2102), DARK D3170(15sec) is OK / night time = 21840(sec)

2007/11/24: vacuum sensor broken, change vacuum sensor(No.44200127 -> No.44200114), punming before observation, change exposure time 240sec due to full moon, DARK D3173(240sec) is OK / night time = 21540(sec)

2007/11/25: DARK D3178(300sec) is OK / night time = 21360(sec)

2007/11/26: DARK D3182(300sec) is OK / night time = 21120(sec)

2007/11/27: Couldn't take good DARK (D3188(300sec) chip2 noisy) / night time = 20940(sec)

2007/11/28: DARK D3199(300sec) is OK / night time = 20760(sec)

2007/11/29: DARK D3203(300sec) is OK / night time = 20520(sec)

2007/11/30: DARK D3206(300sec) is OK. Today was good seeing! / night time = 20340(sec)

TOTAL NIGHT TIME: 706980(sec)

OPERATING RATE: 56.1 %
GB: 3.6 %
LMC: 50.5 %
SMC: 2.0 %
ALERT: 0.0 %
GRB: 0.0 %

Return to index

Previous month Next month